A disk rotates about an axis through its center. Point A is located on its rim and point B is located exactly one fifth of the way from the center toward the rim. What is the ratio of the angular velocity 𝜔A to that of 𝜔B, and the tangential velocity vA to that of vB?


(a) the angular velocity 𝜔A to that of 𝜔B

𝜔A/𝜔B= 


(b) the tangential velocity vA to that of vB

vA

vB = 


1
Expert's answer
2021-10-03T14:13:50-0400

(a)

"ω_A=ω_B \\

\frac{ω_A}{ω_B} = 1"

(b)

"\frac{v_A}{v_B} = \frac{r_Aω_A}{r_Bω_B} = \frac{r_A}{\frac{1}{5}r_A} = 5"